OpenStudy (anonymous):

Vanessa and William are stuck simplifying radical expressions. Vanessa has to simplify the quantity of x to the four thirds power, over x to the five sixths power. William has to simplify the sixteenth root of the quantity of x times x to the third times x to the fourth. Using full sentences describe how to fully simplify Vanessa and William's expressions. Describe if Vanessa and William started with equivalent expressions or if they started with expressions that are not equal.

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

\[\frac{x^m}{x^n} = x^{m-n}\]

OpenStudy (whpalmer4):

\[\sqrt[n]{x} = x^{\frac{1}n}\]

OpenStudy (whpalmer4):

\[x^m*x^n = x^{n+m}\]I think that is all you need...

OpenStudy (math&ing001):

You have to simplify these 2 expressions : \[\frac{ x ^{\frac{ 4 }{ 3 }} }{ x ^{\frac{ 5 }{ 6 }} } and \sqrt[16]{x ^{8}}\] You can use the properties palmer just wrote.

OpenStudy (anonymous):

so @math&ing001 for the the first equation would \[x \frac{ 5 }{ 18 }\] be right

OpenStudy (anonymous):

and for william the final answer would be \[x \frac{ 12 }{ 16 }\]

OpenStudy (math&ing001):

Check again, 4/5 - 5/6 is not equal to 5/18 !

OpenStudy (math&ing001):

Sorry 4/3 - 5/6*

OpenStudy (math&ing001):

And for William's radical expression use this : \[\sqrt[a]{x ^{b}}=x ^{\frac{ b }{ a }}\] Good luck !

OpenStudy (anonymous):

i thought you had to find the gcf

OpenStudy (anonymous):

or least common factoe

OpenStudy (math&ing001):

Yes you do for adding fractions but you made a little mistake, show me the steps in your work, maybe we can find it.

OpenStudy (anonymous):

\[x \frac{ 4 }{ 3 }\div x \frac{ 5 }{ 6}\]

OpenStudy (anonymous):

\[\frac{ 20 }{ 18}-\frac{ 15 }{ 18 }\]

OpenStudy (math&ing001):

See that's the mistake. It's not 20 but 24 \[x ^{\frac{ 4 }{ 3 }-\frac{ 5 }{ 6 }}=x ^{\frac{ 24-15 }{ 18 }}\]

OpenStudy (anonymous):

oh ok so it would be \[\frac{ 9 }{18}=\frac{ 1 }{ 2 }\]

OpenStudy (math&ing001):

Yep !

OpenStudy (anonymous):

ok and what about the other equation

OpenStudy (anonymous):

was my final answer for william was right

OpenStudy (math&ing001):

Nope, see the hint I gave above.

OpenStudy (anonymous):

i thought i did it right because i thought you had to multiply the 3 and the 4

OpenStudy (anonymous):

do i have to switch the number around

OpenStudy (math&ing001):

Use this property : \[\sqrt[a]{x ^{b}}=x ^{\frac{ b }{ a }}\]

OpenStudy (anonymous):

i did already

OpenStudy (math&ing001):

Then how would you get 12/16 ?

OpenStudy (anonymous):

because i thought you was suppose to multiple the 3 and the 4

OpenStudy (anonymous):

am i not suppose to multiply the 3 and 4

OpenStudy (math&ing001):

This is William's expression :\[\sqrt[16]{x ^{8}}\] There's no 3 or 4 in it.

OpenStudy (anonymous):

yes there is inside the radical it has an x*x^3*x^4

OpenStudy (anonymous):

in the problem that im looking at it has that

OpenStudy (math&ing001):

Yeah, I kind of did that part for you. You're not supposed to multiply the 3 and 4, you add all the powers like this x*x^3*x^4=x^(1+3+4)

OpenStudy (anonymous):

ohhh ok so it would be 8/16

OpenStudy (anonymous):

which then equals to 1/2 as well

OpenStudy (math&ing001):

Exactly :)

OpenStudy (airyana1114):

for the first equation is the answer x^1/2 or just 1/2?

OpenStudy (whpalmer4):

\[\frac{x^{4/3}}{x^{5/6}}=x^{(4/3-5/6)}=\] now, how would you get from that to the answer being simply \(1/2\), when you do not know anything about the value of \(x\)?